A peice of unknown metal with a right rectangular prism shape has width of 3.2cm, a length of 17.1cm and height of 4.0cm. It's mass is 1.5kg. Calculate the density of the metal in g/cm3
1
Expert's answer
2020-11-17T10:42:09-0500

m = 1.5 kg = 1500 g

w = 3.2 cm

l = 17.1 cm

h = 4.0 cm

Volume of a right rectangular prism:

"V = l \times w \times h"

"V = 17.1 \times 3.2 \times 4.0 = 218.88 \;cm^3"

The density of the metal "ρ = \frac{m}{V}"

"ρ = \frac{1500}{218.88} = 6.85 \;g/cm^3"